# Test-Time Efficient Pretrained Model Portfolios for Time Series Forecasting

### v1

@@ 1 Introduction
The dominant paradigm in pretrained time series models is the “bigger is better” view, supported by evidence that larger models improve forecast accuracy. Building on this observation, recent work has focused on scaling up both model size and training data to build better zero-shot forecasters (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024a; Shi et al., 2024; Edwards et al., 2025). However, these large, monolithic models come with high training and inference costs, which limits their practical use. To offset the need for ever-larger models, a promising alternative has emerged: allocating extra compute at test time. In NLP and computer vision, this includes generating multiple outputs for a single input and aggregating them—either by sampling from a single model (Sun et al., 2024) or using multiple models (Mavromatis et al., 2024). These examples, which partially inspire our work, belong to a broader class of methods that continue to achieve increasingly successful results across domains. Nevertheless, such strategies have not yet found their way to time series forecasting.
@@
    We show that a portfolio of small pretrained models, each specializing on a subset of the training corpus, can match the accuracy of large monolithic models, while being more compute-efficient. By combining models at test time using ensembling or model selection, our approach achieves accuracy comparable to state-of-the-art pretrained forecasters, while dramatically reducing inference cost. We further observe that portfolios follow similar compute–performance scaling trends as monolithic models.
@@ 2.1 Time series forecasting
Forecasting approaches fall into three broad categories (Benidis et al., 2022; Januschowski et al., 2020). Local models, such as ETS and ARIMA (Hyndman & Athanasopoulos, 2018), train separate models, often fully determined by several parameters ($\sim$10), for each individual time series. Global models, including deep learning methods (Salinas et al., 2020; Rangapuram et al., 2018; Oreshkin et al., 2020; Nie et al., 2023), learn shared parameters ($\sim$100K) across a single dataset of multiple related time series (e.g., energy consumption of different households). More recently, pretrained models such as Chronos (Ansari et al., 2024a), TimesFM (Das et al., 2024), and Moirai (Woo et al., 2024) have adapted ideas from large language models to time series forecasting. These models are trained on large and diverse collections of time series data and aim to generalize to new forecasting tasks without task-specific fine-tuning. With parameter counts typically exceeding 10M, they serve as universal forecasters capable of making accurate zero-shot predictions across a wide range of domains. To reduce inference costs associated with such large models, Ekambaram et al. (2024) proposed Tiny Time Mixers (TTMs)—lightweight models with around 1M parameters, specialized for specific context and forecast settings. Our approach shares the same goal of reducing the test-time cost, but instead of specializing a single small model per setting, we train a portfolio of small models and select or combine them at test time.
@@ 2.2 Model and forecast combination
Traditional model combination methods, such as bagging (Breiman, 1996a) and boosting (Freund & Schapire, 1997), improve prediction accuracy by aggregating the outputs of multiple complementary models. Similar ideas have been applied in time series forecasting under the name forecast combination (Wang et al., 2023b). Forecasts are often combined using weighted averages of individual model predictions. These weights are typically shared across all time series in the dataset and are either set uniformly or in proportion to each model’s validation performance (Pawlikowski & Chorowska, 2020). More flexible strategies learn these weights directly from data, for example using greedy ensemble selection (Caruana et al., 2004; Shchur et al., 2023).
@@
Another related approach for pretrained models is the mixture-of-experts (MoE) framework (Jacobs et al., 1991; Shazeer et al., 2017), where component models and combination rules are jointly trained. Although this is shown to be effective for reducing inference time, it limits flexibility. In contrast, our approach trains the component models in the portfolio independently, and allows the use of different fusion strategies at test-time. Notably, this flexibility enables the fusion rule to be adapted according to available computational resources during inference. Additionally, our approach enhances interpretability during the prediction by clarifying the contribution of each portfolio member, whose training procedures are transparent.
@@ 3.4 Forecast combination at test-time
Given a new time series dataset that was not seen by the portfolio members during training, we can utilize the models through one of two strategies: model selection or ensembling. In the first approach, we simply select the single best-performing model based on its accuracy on a validation set. In the second approach, we combine the predictions from multiple models using weighted averaging $\hat{y}_{\mathrm{ens}}=\sum_{m=1}^{M}w_{m}\cdot\hat{y}_{m}$, where the weights are determined using the ensemble selection algorithm (Caruana et al., 2004). For completeness, we provide the full description of the ensemble selection algorithm in Appendix D. Both strategies depend on validation sets produced through time series cross-validation (Hyndman & Athanasopoulos, 2018) to rank or weight the predictions in the model portfolio.
@@ 4 Experiments and Results
For each model size, we first train a generalist using the original Chronos-Bolt corpus and a 200K-gradient-step training procedure following the reference implementations Ansari et al. (2024b), and incorporating mixup and synthetic data as specified in Ansari et al. (2024a). We then construct a portfolio of specialists by further fine-tuning (i.e., post-training) the generalist weights on different subsets of the training corpus for 1K gradient steps.
In exploratory experiments, we also tested 3K, 10K, and 20K post-training steps, but observed no significant performance differences. As a result, we did not tune this variable further or include it as an experimental factor, and we use 1K gradient steps in our main experiments. This post-training phase amounts to only 0.5% of the generalist’s training time, making our method of constructing specialists extremely lightweight. Each portfolio includes the generalist and its associated specialists.
To evaluate a portfolio on a target task, we hold out the last $H$ observations from the training set to create a validation window. All models in the portfolio generate forecasts for this window. We then either select the model with the lowest loss on the validation set, or fit a weighted ensemble using the greedy ensemble selection algorithm (Caruana et al., 2004) implemented in AutoGluon (Shchur et al., 2023). The selected model or ensemble is then applied to the test set to produce final forecasts.
Most of our results are reported on Chronos Benchmark II [^2] https://huggingface.co/spaces/autogluon/fev-leaderboard (BM2). Importantly, all datasets in the benchmark are zero-shot examples for our models (these datasets, including their training partitions, were not used in training of Chroma portfolios). Following the benchmark, we report the probabilistic forecasting results using the weighted quantile loss (WQL) and point forecasting results using the mean absolute scaled error (MASE). Finally, we also evaluate the larger Chroma variants on GIFT-Eval [^3] https://huggingface.co/spaces/Salesforce/GIFT-Eval (Aksu et al., 2024). For both benchmarks, we report aggregate performance using the geometric mean of relative errors. Specifically, for each dataset, we divide the model’s error by that of a baseline (Seasonal Naive unless stated otherwise), and then compute the geometric mean across datasets (Fleming & Wallace, 1986). Further details on the experiment setup are given in Appendix B.
@@ 4.1 Performance and test-time efficiency
We report overall results for Chroma on BM2 in Figure 2 (left), comparing the performance of 4m and tiny (9m) Chroma portfolios of frequency specialists to the state of the art in pretrained time series models. We find that, despite having as little as 4M active parameters at test time, the Chroma portfolio performs comparably to much larger monoliths such as Moirai-1.1 Large (311M parameters) (Woo et al., 2024), TimesFM-2.0 (500M parameters) (Das et al., 2024) or Chronos-Bolt Base (205M parameters).
We also report results on the GIFT-Eval benchmark, in Figure 2 (right), including only pretrained architectures and statistical baselines for comparison. In rolling window evaluation tasks, we perform model selection or ensembling on the Chroma model portfolio on the first evaluation window alone, and infer with this fixed selection across the subsequent windows. In this benchmark, too, Chroma’s aggregated probabilistic forecasting performance over 97 tasks is close to some much larger pretrained architectures, including TabPFN-TS (Hoo et al., 2025) which leverages significant test-time computation for in-context learning with a tabular foundation model. Note that we report results for TTM (Ekambaram et al., 2024) in zero-shot mode only. Moreover, in order not to misrepresent the performance of TTM, which specializes in high-frequency time series, we omit TTM’s results on BM2 which primarily contains low-frequency time series datasets. Further details on benchmark evaluations, including point forecasting performance, are given in Appendix G.
@@
We observe that building time series model portfolios by partitioning datasets by frequency is preferable to partitioning them by application domain, as this results in $\sim 5\%$ decrease in overall MASE and $\sim 4\%$ in WQL, across all model sizes considered (Table 1). This is consistent with earlier work such as TTM (Ekambaram et al., 2024) who used frequency to partition datasets. When comparing the use of ensemble vs. model selection, although ensemble selection leads to a slight improvement across model sizes and metrics, we find the evidence for this is not as strong as our previous conclusion. Therefore, while using ensembling to combine Chroma models can result in improvements in accuracy, this difference may not be strong enough to justify the added computational costs of using multiple models for inference. Finally, we observe that simply combining multiple generalists trained using the same large data set does not result in significant performance gains over just using a single generalist. We hypothesize this is due to variance being a negligible factor in the test error committed by models pretrained on large corpora—an observation that can be supported by Lin et al. (2024), who argues that current LLMs remain underfitted owing to the relatively small number of training epochs. We discuss this observation further in Appendix F.
@@
Having observed that our conclusions hold across different model sizes, we turn our attention to quantifying how Chroma scales. Similar to recent works (Hoffmann et al., 2022; Kaplan et al., 2020; Shi et al., 2024; Edwards et al., 2025), we carry out ordinary least squares fits to the test error of Chroma (on BM2) and the number of active parameters at inference time. We present results for model selection on the Chroma portfolio in Figure 3. We observe that despite individually training on much smaller datasets, the overall model portfolio’s performance scales at rates comparable to the individual generalist models. In further experiments, we found that our training setup of post-training on generalists to obtain the portfolio is essential to this conclusion (see, Appendix G). We therefore conclude that, provided datasets can be proportionally scaled, our approach can be scaled up to larger model sizes to obtain even better performance.
@@ 4.2 Further experiments
Second, we observe that the benefit of specialist portfolios comes primarily from the diversity they introduce. However, this diversity is only useful if we combine models intelligently at test time. In particular, our ensembles do not work simply by reducing variance through averaging, as might be expected in settings where randomness in training dominates model differences. We confirm this in ablation studies: both simple averaging and performance-weighted averaging (cf. Pawlikowski & Chorowska, 2020) lead to significant performance drops when applied to our pretrained portfolios. This contrasts with much of the prior forecasting literature, where ensembles of task-specific models are often effective even with naive combination strategies (Wang et al., 2023b).
@@ 5 Conclusion
In this work, we introduced Chroma: a portfolio of small, pretrained time series forecasters. Chroma consists of specialist models, each fine-tuned on disjoint subsets of the training data from a generalist pretrained model. When model selection or ensembling is applied at test time, Chroma achieves competitive performance on forecasting benchmarks from Ansari et al. (2024a) and Aksu et al. (2024), matching the accuracy of much larger monolithic models. Our experiments demonstrate the test-time efficiency of Chroma, as well as its scalability. We believe that the methodology behind Chroma opens a promising new direction for enhancing the accuracy of pretrained forecasting models, especially under limited compute budgets. Furthermore, while our focus is on time series forecasting, the underlying principles may also extend to other domains such as natural language processing and computer vision. Our strategy offers an alternative to approaches like Best-of-$N$ which samples from a single base model. Instead, it proposes an inference-aware portfolio formation strategy: train a single generalist base model, fine-tune it into a portfolio of specialized experts, and apply model selection at test time.
@@
We focused on test-time efficiency because inference often dominates the computational budget in deployed forecasting systems. To that end, we showed that small portfolios of pretrained models can match the performance of much larger models while requiring substantially less compute at test time. Although we did not study the impact of the pretraining phase in detail, our design choices follow standard practice (see Appendix B. Also, our theoretical observations in Appendix F provides a bias-variance perspective that helps explain the observed gains and suggests implications for the role of pretraining. Understanding how the number of training iterations and other pretraining hyperparameters interact with our bias–variance explanation would be a valuable direction for future research. Moreover, while we hand-crafted the specialist portfolios using known dataset features (e.g., domain or frequency), future work could explore more principled approaches. For instance, foundation model specialists can be formed by incorporating boosting-style objectives (Freund & Schapire, 1997). Finally, this work focused on weighted ensembling; future research could explore more sophisticated methods such as stacked ensembling (Breiman, 1996b) that can lead to higher accuracy.
@@ Appendix A Additional Literature
Prior work shows that ensembles of neural networks outperform single networks and provide better uncertainty quantification. Specifically, it has been reported that training the same model architecture from different random initializations is an effective strategy for sufficiently diverse portfolios (Lee et al., 2015), and deep ensembles utilize this idea (Lakshminarayanan et al., 2017; Fort et al., 2019; Zaidi et al., 2021). At the foundation-model scale, however, such work is less common. Notably, Sun et al. (2022) builds LLM ensembles by varying random seeds, while Wang et al. (2023a) uses low-rank adapters for resource efficiency. We adopt a similar approach while forming our generalist portfolios and likewise observe gains over single models. However, we further show that larger improvements can be achieved with specialist portfolios, where models are fine-tuned on targeted subsets of the data distribution. As discussed in Appendix F, in the foundation-model regime, variance reduction from ensembling is less decisive than improved coverage of heterogeneous data.
In time-series context, we also find that partitioning by frequency to induce specialization across portfolio members performs particularly well. This aligns with a broader line of work that exploits frequency-based representation and basis expansions for time-series forecasting. Namely, Darlow et al. (2024) proposes a foundation model that predicts coefficients of sinusoidal bases for functional decomposition, and Kim et al. (2024) represents time series directly in the frequency space. As opposed to those methods aiming to learn a single, monolithic model, our methodology is focused on specialization, i.e., training multiple experts on frequency-based clusters and aggregating their predictions. In classical statistical forecasting, beyond our baselines such as Auto-ETS and Auto-ARIMA, there are also recent methods such as FFORMA (Montero-Manso et al., 2020), which learns to predict ensemble weights over a pool of univariate forecasters using time-series features, and SCUM (Petropoulos & Svetunkov, 2020), which aggregates established statistical methods via median-based fusion.

### v2

@@ 1 Introduction
The dominant paradigm in pretrained time series models is the “bigger is better” view, supported by evidence that larger models improve forecast accuracy. Building on this observation, recent work has focused on scaling up both model size and training data to build better zero-shot forecasters (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024a; Shi et al., 2024a; Edwards et al., 2025). However, these large, monolithic models come with high training and inference costs, which limits their practical use. To offset the need for ever-larger models, a promising alternative has emerged: allocating extra compute at test time. In NLP and computer vision, this includes generating multiple outputs for a single input and aggregating them—either by sampling from a single model (Sun et al., 2024) or using multiple models (Mavromatis et al., 2024). These examples, which partially inspire our work, belong to a broader class of methods that continue to achieve increasingly successful results across domains. Nevertheless, such strategies have not yet found their way to time series forecasting.
@@
    We show that a portfolio of small pretrained models (based on encoder-decoder Chronos-Bolt Ansari et al. (2024b) architecture), each specializing on a subset of the training corpus, can match the accuracy of large monolithic models, while being more compute-efficient. By combining models at test time using ensembling or model selection, our approach achieves accuracy comparable to state-of-the-art pretrained forecasters, while dramatically reducing inference cost. We further observe that portfolios follow similar compute–performance scaling trends as monolithic models.
@@ 2.1 Time series forecasting
Forecasting approaches fall into three broad categories (Benidis et al., 2022; Januschowski et al., 2020). Local models, such as ETS and ARIMA (Hyndman and Athanasopoulos, 2018), train separate models, often fully determined by several parameters ($\sim$10), for each individual time series. Global models, including deep learning methods (Salinas et al., 2020; Rangapuram et al., 2018; Oreshkin et al., 2020; Nie et al., 2023), learn shared parameters ($\sim$100K) across a single dataset of multiple related time series (e.g., energy consumption of different households). More recently, pretrained models such as Chronos (Ansari et al., 2024a), TimesFM (Das et al., 2024), and Moirai (Woo et al., 2024) have adapted ideas from large language models to time series forecasting. These models are trained on large and diverse collections of time series data and aim to generalize to new forecasting tasks without task-specific fine-tuning. With parameter counts typically exceeding 10M, they serve as universal forecasters capable of making accurate zero-shot predictions across a wide range of domains. To reduce inference costs associated with such large models, Ekambaram et al. (2024) proposed Tiny Time Mixers (TTMs)—lightweight models with around 1M parameters, specialized for specific context and forecast settings. Our approach shares the same goal of reducing the test-time cost, but instead of specializing a single small model per setting, we train a portfolio of small models and select or combine them at test time.
@@ 2.2 Model and forecast combination
Traditional model combination methods, such as bagging (Breiman, 1996a) and boosting (Freund and Schapire, 1997), improve prediction accuracy by aggregating the outputs of multiple complementary models. Similar ideas have been applied in time series forecasting under the name forecast combination (Wang et al., 2023b). Forecasts are often combined using weighted averages of individual model predictions. These weights are typically shared across all time series in the dataset and are either set uniformly or in proportion to each model’s validation performance (Pawlikowski and Chorowska, 2020). More flexible strategies learn these weights directly from data, for example using greedy ensemble selection (Caruana et al., 2004; Shchur et al., 2023).
@@
Another related approach for pretrained models is the mixture-of-experts (MoE) framework (Jacobs et al., 1991; Shazeer et al., 2017), where component models and combination rules are jointly trained. This approach is shown to be effective in time-series forecasting tasks as well (Shi et al., 2024b; Liu et al., 2024). Although these works showed that MoE can help reducing inference time and cost, it limits flexibility. In contrast, our approach trains the component models in the portfolio independently, and allows the use of different fusion strategies at test-time. Notably, this flexibility enables the fusion rule to be adapted according to available computational resources during inference. Additionally, our approach enhances interpretability during the prediction by clarifying the contribution of each portfolio member, whose training procedures are transparent.
@@ 3.4 Forecast combination at test-time
Given a new time series dataset that was not seen by the portfolio members during training, we can utilize the models through one of two strategies: model selection or ensembling. In the first approach, we simply select the single best-performing model based on its accuracy on a validation set. In the second approach, we combine the predictions from multiple models using weighted averaging $\hat{y}_{\mathrm{ens}}=\sum_{m=1}^{M}w_{m}\cdot\hat{y}_{m}$, where the weights are determined using the ensemble selection algorithm (Caruana et al., 2004). For completeness, we provide the full description of the ensemble selection algorithm in Appendix D. Both strategies depend on validation sets produced through time series cross-validation (Hyndman and Athanasopoulos, 2018) to rank or weight the predictions in the model portfolio.
@@ 4 Experiments and Results
For each model size, we first train a generalist using the original Chronos-Bolt corpus and a 200K-gradient-step training procedure following the reference implementations Ansari et al. (2024b), and incorporating mixup and synthetic data as specified in Ansari et al. (2024a). We then construct a portfolio of specialists by further fine-tuning (i.e., post-training) the generalist weights on different subsets of the training corpus for 1K gradient steps. In exploratory experiments, we also tested 3K, 10K, and 20K post-training steps, but observed no significant performance differences. As a result, we did not tune this variable further or include it as an experimental factor, and we use 1K gradient steps in our main experiments. This post-training phase amounts to only 0.5% of the generalist’s training time, making our method of constructing specialists extremely lightweight. Each portfolio includes the generalist and its associated specialists.
To evaluate a portfolio on a target task, we hold out the last $H$ observations from the training set to create a validation window. All models in the portfolio generate forecasts for this window. We then either select the model with the lowest loss on the validation set, or fit a weighted ensemble using the greedy ensemble selection algorithm (Caruana et al., 2004) implemented in AutoGluon (Shchur et al., 2023). These are based on rolling-window evaluations (also referred to as backtesting or out-of-time folds) and can be performed even with a single time series. It can also be performed by increasing the number of evaluation windows. Namely, instead of using a single window for selecting which specialists provide the best forecasts, one could rely on multiple. The selected model or ensemble is then applied to the test set to produce final forecasts. Note that we use the same methodology throughout all experiments.
Most of our results are reported on Chronos Benchmark II [^2] https://huggingface.co/spaces/autogluon/fev-leaderboard (BM2). Importantly, all datasets in the benchmark are zero-shot examples for our models (these datasets, including their training partitions, were not used in training of Chroma portfolios). Following the benchmark, we report the probabilistic forecasting results using the weighted quantile loss (WQL) and point forecasting results using the mean absolute scaled error (MASE). Finally, we also evaluate the larger Chroma variants on GIFT-Eval [^3] https://huggingface.co/spaces/Salesforce/GIFT-Eval (Aksu et al., 2024). For both benchmarks, we report aggregate performance using the geometric mean of relative errors. Specifically, for each dataset, we divide the model’s error by that of a baseline (Seasonal Naive unless stated otherwise), and then compute the geometric mean across datasets (Fleming and Wallace, 1986). Further details on the experiment setup are given in Appendix B.
@@ 4.1 Performance and test-time efficiency
We report overall results for Chroma on BM2 in Figure 2 (left), comparing the performance of 4m and tiny (9m) Chroma portfolios of frequency specialists to the state of the art in pretrained time series models. We find that, despite having as little as 4M active parameters at test time, the Chroma portfolio under best model selection performs comparably to much larger monoliths such as Moirai-1.1 Large (311M parameters) (Woo et al., 2024), TimesFM-2.0 (500M parameters) (Das et al., 2024) or Chronos-Bolt Base (205M parameters).
We also report results on the GIFT-Eval benchmark, in Figure 2 (right), including only pretrained architectures and statistical baselines for comparison. In rolling window evaluation tasks, we perform model selection or ensembling on the Chroma model portfolio on the first evaluation window alone, and infer with this fixed selection across the subsequent windows. That is, the ensemble was fixed in the training window of the first rolling window, and was not refit in subsequent windows. In this benchmark, too, Chroma’s aggregated probabilistic forecasting performance over 97 tasks is close to some much larger pretrained architectures, including TabPFN-TS (Hoo et al., 2025) which leverages significant test-time computation for in-context learning with a tabular foundation model. Note that we report results for TTM (Ekambaram et al., 2024) in zero-shot mode only. Moreover, in order not to misrepresent the performance of TTM, which specializes in high-frequency time series, we omit TTM’s results on BM2 which primarily contains low-frequency time series datasets. Further details on benchmark evaluations, including point forecasting performance, are given in Appendix G.
@@
We observe that building time series model portfolios by partitioning datasets by frequency is preferable to partitioning them by application domain, as this results in $\sim 5\%$ decrease in overall MASE and $\sim 4\%$ in WQL, across all model sizes considered (Table 1). This is consistent with earlier work such as TTM (Ekambaram et al., 2024) who used frequency to partition datasets. When comparing the use of ensemble vs. model selection, although ensemble selection leads to a slight improvement across model sizes and metrics, we find the evidence for this is not as strong as our previous conclusion. Therefore, while using ensembling to combine Chroma models can result in improvements in accuracy, this difference may not be strong enough to justify the added computational costs of using multiple models for inference.
Finally, we observe that simply combining multiple generalists trained using the same large data set does not result in significant performance gains over just using a single generalist. We hypothesize this is due to variance being a negligible factor in the test error committed by models pretrained on large corpora—an observation that can be supported by Lin et al. (2024), who argues that current LLMs remain underfitted owing to the relatively small number of training epochs. We discuss this observation further in Appendix F. Specifically, we show that in the regime where Chroma models are trained, bias component of the generalization error dominates the error due to variance. Since specialists reduce bias within a specific part of the overall data distribution, intelligent model combination or ensembling at test-time can reduce the overall bias of the portfolio. This is as opposed to the traditional ensembling approaches like bagging that tries to reduce error by decreasing variance.
@@
Having observed that our conclusions hold across different model sizes, we turn our attention to quantifying how Chroma scales. Similar to recent works (Hoffmann et al., 2022; Kaplan et al., 2020; Shi et al., 2024a; Edwards et al., 2025), we carry out ordinary least squares fits to the test error of Chroma (on BM2) and the number of active parameters at inference time. We present results for model selection on the Chroma portfolio in Figure 3. We observe that despite individually training on much smaller datasets, the overall model portfolio’s performance scales at rates comparable to the individual generalist models. In further experiments, we found that our training setup of post-training on generalists to obtain the portfolio is essential to this conclusion (see, Appendix G). We therefore conclude that, provided datasets can be proportionally scaled, our approach can be scaled up to larger model sizes to obtain even better performance.
@@ 4.2 Further experiments
Second, we observe that the benefit of specialist portfolios comes primarily from the diversity they introduce. However, this diversity is only useful if we combine models intelligently at test time. In particular, our ensembles do not work simply by reducing variance through averaging, as might be expected in settings where randomness in training dominates model differences. We confirm this in ablation studies: both simple averaging and performance-weighted averaging (cf. Pawlikowski and Chorowska, 2020) lead to significant performance drops when applied to our pretrained portfolios. This contrasts with much of the prior forecasting literature, where ensembles of task-specific models are often effective even with naive combination strategies (Wang et al., 2023b).
@@ 5 Conclusion
In this work, we introduced Chroma: a portfolio of small, pretrained time series forecasters. Chroma consists of specialist models, each fine-tuned on disjoint subsets of the training data from a generalist pretrained model. When model selection or ensembling is applied at test time, Chroma achieves competitive performance on forecasting benchmarks from Ansari et al. (2024a) and Aksu et al. (2024), matching the accuracy of much larger monolithic models. Our experiments demonstrate the test-time efficiency of Chroma, as well as its scalability. We believe that the methodology behind Chroma opens a promising new direction for enhancing the accuracy of pretrained forecasting models, especially under limited compute budgets. Furthermore, while our focus is on time series forecasting, the underlying principles may also extend to other domains such as natural language processing and computer vision. Our strategy offers an alternative to approaches like Best-of-$N$ which samples from a single base model. Instead, it proposes an inference-aware portfolio formation strategy: train a single generalist base model, fine-tune it into a portfolio of specialized experts, and apply model selection at test time. Moreover, we focused on test-time efficiency because inference often dominates the computational budget in deployed forecasting systems.
@@
Although we did not study the impact of the pretraining phase in detail, our design choices follow standard practice: T5-architecture based Chronos-Bolt model with straightforward modifications, see Appendix B. Also, our theoretical observations in Appendix F provides a bias-variance perspective that helps explain the observed gains and suggests implications for the role of pretraining. Understanding how the number of training iterations and other pretraining hyperparameters interact with our conclusions would be a valuable direction for future research. Similarly, our experiments were conducted with encoder-decoder transformer architectures. Future work can investigate portfolios of transformers based on other types of architectures. Moreover, while we hand-crafted the specialist portfolios using known dataset features (e.g., domain or frequency), future work could explore more principled approaches. For instance, foundation model specialists can be formed by incorporating boosting-style objectives (Freund and Schapire, 1997). Finally, this work focused on weighted ensembling; future research could explore more sophisticated methods such as stacked ensembling (Breiman, 1996b) that can lead to higher accuracy.
@@ Appendix A Additional Literature
Prior work shows that ensembles of neural networks outperform single networks and provide better uncertainty quantification. Specifically, it has been reported that training the same model architecture from different random initializations is an effective strategy for sufficiently diverse portfolios Lee et al. (2015), and deep ensembles utilize this idea Lakshminarayanan et al. (2017); Fort et al. (2019); Zaidi et al. (2021). At the foundation-model scale, however, such work is less common. Notably, Sun et al. (2022) builds LLM ensembles by varying random seeds, while Wang et al. (2023a) uses low-rank adapters for resource efficiency. We adopt a similar approach while forming our generalist portfolios and likewise observe gains over single models. However, we further show that larger improvements can be achieved with specialist portfolios, where models are fine-tuned on targeted subsets of the data distribution. As discussed in Appendix F, in the foundation-model regime, variance reduction from ensembling is less decisive than improved coverage of heterogeneous data.
In time-series context, we also find that partitioning by frequency to induce specialization across portfolio members performs particularly well. This aligns with a broader line of work that exploits frequency-based representation and basis expansions for time-series forecasting. Namely, Darlow et al. (2024) proposes a foundation model that predicts coefficients of sinusoidal bases for functional decomposition, and Kim et al. (2024) represents time series directly in the frequency space. As opposed to those methods aiming to learn a single, monolithic model, our methodology is focused on specialization, i.e., training multiple experts on frequency-based clusters and aggregating their predictions. In classical statistical forecasting, beyond our baselines such as Auto-ETS and Auto-ARIMA, there are also recent methods such as FFORMA (Montero-Manso et al., 2020), which learns to predict ensemble weights over a pool of univariate forecasters using time-series features, and SCUM (Petropoulos and Svetunkov, 2020), which aggregates established statistical methods via median-based fusion.